OpenStudy (anonymous):

Can someone private message me? I need a TON of algebra 2 help, and I'd really like to explain the problem in a more private manner.

OpenStudy (anonymous):

what you need?

OpenStudy (anonymous):

Pm me? Dx

OpenStudy (anonymous):

Fan me so I can message you back.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

The design of a digital box camera maximizes the volume while keeping the sum of the dimensions at 6 inches. If the length must be 1.5 times the height, what should each dimension be? Hint: Let x represent one of the dimensions, and then define the other dimensions in terms of x. (2 points)

OpenStudy (anonymous):

the lenght is \(1.5\) times the height so if the height is \(x\) the lenght is \(1.5x\)

OpenStudy (anonymous):

Okay. Gotcha.

OpenStudy (anonymous):

oh no it doesn't say that does it, scratch that lets make the width \(y\) so \[x+1.5x+y=6\] or \[2.5x+y=6\]

OpenStudy (anonymous):

L+w+h right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay I see that.

OpenStudy (anonymous):

now we want to maximize the volume \[V=LWH=x\times 1.5x\times y\] we need to solve for \(y\)

OpenStudy (anonymous):

\[2.5x+y=6\] so \[y=6-2.5x\]

OpenStudy (anonymous):

now the volume is \[V(x)=1.5x^2(6-2.5x)\]

OpenStudy (anonymous):

is this a calculus class?

OpenStudy (anonymous):

Algebra 2..

OpenStudy (anonymous):

hmm i am stuck without calculus because you have a cubic funcition the answer is H = 1.6 L = 2.4 W = 2.0

OpenStudy (anonymous):

Okay. Thank you.. I've been stuck on that one for awhile. I'm working on another, if you give me a second I'll post the next.

OpenStudy (anonymous):

ok really i have no idea how you do that without calc

OpenStudy (anonymous):

I'm seriously at a loss of words, some of this is beyond me. I don't even take calc next, just pre calc.

OpenStudy (anonymous):

ok lets do one more them maybe some tomorrow you have time tomorrow?

OpenStudy (anonymous):

Use Pascal’s triangle to expand the binomial. (d – 5y)6 (2 points)

OpenStudy (anonymous):

almost all day tomorrow.

OpenStudy (anonymous):

ok i see it says "two points" so perhaps we can just get the answer, is that good enough?

OpenStudy (anonymous):

This one I have to work out. If you just explain it to me I will do all of the work.

OpenStudy (anonymous):

yikes ok first the answer, it is \[d^6-30 d^5 y+375 d^4 y^2-2500 d^3 y^3+9375 d^2 y^4-18750 d y^5+15625 y^6\]

OpenStudy (anonymous):

now the working out we need the sixth level of pascals triangle, the one that begins with 1 6

OpenStudy (anonymous):

i don't know if you have a picture of pascal's triangle, you can google it, but the sixth level has the numbers \[1,5,15,20,15,6,1\]

OpenStudy (anonymous):

those are the coefficents we need then \[(a+b)^6=a^6+6a^5b+15a^4b^2+20a^3b^3+15a^2b^4+6ab^5+b^6\]

OpenStudy (anonymous):

Okay. Ugh.

OpenStudy (anonymous):

where you see an \(a\) put a \(d\) where you see a \(b\) put \(-5d\)

OpenStudy (anonymous):

you can see why the numbers at the end get so big

OpenStudy (anonymous):

Okay.. So.. one sec.

OpenStudy (anonymous):

in any case if you do it right you get \[d^6-30 d^5 y+375 d^4 y^2-2500 d^3 y^3+9375 d^2 y^4-18750 d y^5+15625 y^6\]

OpenStudy (anonymous):

I'm going to have faith in you. Thank you..

OpenStudy (anonymous):

actually i did not do it by hand, it is donkey work i did it here http://www.wolframalpha.com/input/?i=%28d-5y%29^6

OpenStudy (anonymous):

x4 – 41x2 = –400 I have to find the real and complex solutions for this, I know it's not too hard.

OpenStudy (anonymous):

let me refresh i got the dreaded quesiton marks

OpenStudy (anonymous):

no this one i s going to be easy

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

\[x^4-41 x^2+400 = 0\] is a start then this one actually factors as \[(x^2-25)(x^2-16)=0\]

OpenStudy (anonymous):

factor again , get \[(x+5)(x-5)(x+4)(x-4)=0\] set each factor equal to zero and solve in your head, get \[x=5,x=-5,x=4,x=-4\]

OpenStudy (anonymous):

snappy one more?`

OpenStudy (anonymous):

That was easy, okay. One moment.

OpenStudy (anonymous):

If 2 + square root 3 is a polynomial root, name another root of the polynomial, and explain how you know it must also be a root.

OpenStudy (anonymous):

I know how to find roots, but when I already have one I don't know how to plug it in?

OpenStudy (anonymous):

it is the "conjugate' \(2-\sqrt3\)

OpenStudy (anonymous):

that is all, they come in conjugate pairs if the coefficients are integers then it must be \(2-\sqrt3\) because \[(2+\sqrt3)(2-\sqrt3)=4-3=1\] an integer

OpenStudy (anonymous):

Okay! Just the opposite!

OpenStudy (anonymous):

you don't have to "find" it because you are told that one is \(2+\sqrt3\) the other is the conjugate that is all not "opposite" but "conjugate"

OpenStudy (anonymous):

Conjugate, got it. (: Should probably sound smarter than I claim to be. Thank you so much.

OpenStudy (anonymous):

yw get some rest